OpenStudy (anonymous):

Which point is NOT a solution of x + 2 y is less than or equal to 8 A. (0, 4) B. (-2, 3) C. (4, 4) D. (6, 0)

OpenStudy (anonymous):

ca someone answer this question plzzzz:)

OpenStudy (campbell_st):

so the equation is \[x + 2y \le 8\] substitute each point e.g (0, 4) x = 0 and y = 4 \[0 + 2 \times 8 \le 8\] that is true, so the point is a solution... substitute each pair to find the point that is NOT a solution hope it helps